OpenStudy (anonymous):

What is the geometric mean of 3 and 21? A. 63 B. 31.563 C. 3 radical 7 D. 7 radical 3

OpenStudy (anonymous):

Geometric mean of 3 and 21 is: √ (3×21) = 7.937253933 3×√7 = 3×2.645751311 = 7.937253933

OpenStudy (anonymous):

Thank u help alot

OpenStudy (anonymous):

No problem :)